# What sets the metallicity of Ultra-Faint Dwarfs?

### v1

@@ 1 Introduction
Currently, there is no established physical interpretation of the metallicities of UFD galaxies and their scatter in terms of the physical processes shaping the evolution of these galaxies. In this study, we attempt to construct such an interpretation using a combination of the IGM metallicity distributions derived from several state-of-the-art cosmological simulations of galaxy formation and a semi-analytical model of galaxy formation based on a large number of mass assembly histories (MAHs) of UFD progenitors residing in the Milky Way-sized halos at $z=0$. Specifically, we explore a range of models with different assumptions in modeling the accreted IGM metallicity and different outflow mass loading factors. We compare the luminosity–metallicity relation of the model galaxies and their stellar metallicity distribution functions (MDFs) to existing measurements for observed UFDs with the goal of gauging the relative importance of external IGM enrichment and outflows in setting UFD metallicities.
@@ 2.1 Model of UFD evolution
Our model for the evolution of ultra-faint galaxies is based on the halo evolution tracks extracted from the Caterpillar [^1] https://www.caterpillarproject.org suite of cosmological zoom-in $N$-body simulations of Milky Way-mass ($0.5\lesssim M/10^{12}\,M_{\odot}\lesssim 2$) halos (Griffen et al., 2016). Specifically, we use the $32$ LX14 Caterpillar simulations with a mass resolution of $3\times 10^{4}M_{\odot}$, which have been shown to be sufficient to reliably model subhalos down to masses of $\gtrsim 10^{7}\,M_{\odot}$ (Pham et al., 2023) corresponding to $V$-band magnitudes of $M_{V}\gtrsim 0$ (Manwadkar & Kravtsov, 2022). Combining these suites provides a population of several thousand halos in MW-like environments, which we use in our study.
@@
where $Z_{\rm IGM}$ is the characteristic metallicity of heavy elements accreted by the galaxy from the IGM, $y_{\rm Z}$ is the yield of heavy elements per unit star formation rate produced by young massive stars and dispersed by supernovae and AGB stars, $\mathcal{R}$ is the fraction of mass lost by stars due to winds and supernovae, assumed to be instantaneous in the model, $Z_{\rm g}=M_{Z,\rm g}/M_{\rm g}$ is the mass fraction of heavy elements in the ISM, $\eta$ is the wind mass loading factor (defined as the ratio of the gas mass outflow rate to the star formation rate) and $\zeta_{\rm w}$ is the wind metallicity enhancement factor $\zeta_{\rm w}=Z_{\rm w}/Z_{\rm g}$ where $Z_{\rm w}$ is the metallicity of the outflow. Finally, $\dot{\mathcal{M}}_{\star}$ is the instantaneous star formation rate of the model galaxy.
@@
In cosmological simulations such as FIRE (Muratov et al., 2015), FIRE-2 (Pandya et al., 2021), EAGLE (Mitchell et al., 2020), and Illustris TNG50 (Nelson et al., 2019b), a clear trend of increasing average $\eta$ with decreasing stellar mass is present. However, these analyses do not probe galaxies with stellar mass as low as UFDs $(M_{\star}\lesssim 10^{5.5}M_{\odot})$, so the extent to which this relation can be extrapolated to galaxies of stellar mass $M_{\star}<10^{6}\,M_{\odot}$ is unclear. Furthermore, the mass loading factor of galaxies at high redshifts where UFDs are expected to form most of their stars is also largely unknown.
@@
The fiducial values of Manwadkar & Kravtsov (2022) adopted in this work are $\eta_{\rm norm}=1.8$, $\eta_{p}=-0.45$, $\eta_{c}=4.5$, $M_{\star,10}\equiv M_{\star}/10^{10}M_{\odot}$, and $\eta_{\rm max}=2000$. These values were chosen to be close to those in the FIRE and FIRE-2 simulations for galaxies of $10^{6}\lesssim M_{\star}/M_{\odot}\lesssim 10^{9}$ (Muratov et al., 2015; Pandya et al., 2021), while at smaller and larger masses the power law relation $\eta\propto M_{\star,10}^{\eta_{p}}$ is modified to reproduce the observed stellar mass to halo mass relation (SHMR) and stellar mass–metallicity relations, as detailed in Manwadkar & Kravtsov (2022). The subtraction of $\eta_{\rm c}$ results in the rapid decrease of $\eta$ for $M_{\star}\gtrsim 10^{10}\,M_{\odot}$, while the maximum value of $\eta_{\rm max}=2000$ was chosen to account for the expectation in the UFD regime the decreasing stellar mass-to-halo mass ratio results in less efficient driving in outflows and to prevent unlimited increase of $\eta$ at the very low stellar masses.
The values of other parameters are set to the fiducial values adopted in Manwadkar & Kravtsov (2022): $y_{Z}=0.06$ for the Chabrier (2003) IMF (e.g., Vincenzo et al., 2016), $\mathcal{R}=0.34$, $\zeta_{\rm w}=1$. The expected value and scaling of $\zeta_{\rm w}$ with system mass are currently not known. Manwadkar & Kravtsov (2022) showed that a good match to the observed luminosity–metallicity relation of bright dwarf galaxies can be obtained with $\zeta_{\rm w}=1$, as long as a mass-dependent scaling of $\eta$ is adopted. Results of the FIRE galaxy formation simulations are also consistent with $\zeta_{\rm w}\approx 1$ independent of galaxy mass (Muratov et al., 2017; Pandya et al., 2021), albeit with a sizable scatter.
@@ 2.2 Galaxy formation simulations
We consider the properties of these simulated MW-like environments in the redshift range $105<z<$ 10, corresponding to the redshift range of the publicly available FIRE-2 high redshift suite. We found that the metallicity of the IGM at $z\approx 8-10$ has the most effect on the stellar metallicities of model UFDs within this range (see Section 3.1). We use the $z=10$ metallicity distribution to model IGM metallicity in the GRUMPY model at $z>10$ – an overestimate of the metallicity of accreted gas, but a conservative choice that strengthens our conclusions below. Similarly, we extend the $z=5$ metallicity distribution to $z=0$ in the GRUMPY model, where the choice of $Z_{\rm IGM}$ has little impact on model UFD metallicities because UFDs are no longer star forming.
@@ 2.2.1 FIRE-2 simulations
We consider simulation outputs in the redshift range $105<z<$ 10, the range publicly available in both the core m12 models (Wetzel et al., 2016; Garrison-Kimmel et al., 2017, 2019; Hopkins et al., 2018; Samuel et al., 2020) and the high redshift simulations of smaller-mass halos (e.g., Ma et al., 2018, 2019, 2020), suitable to probe the environments of UFDs prior to cosmic reionization when most of their stars are expected to form. All FIRE-2 simulations analyzed here assume an IGM gas metallicity floor of $\mathrm{[Fe/H]}\simeq-3.9$.
2.2.2 Semenov et al. 2024 simulation of a Milky Way progenitor
We also study a simulation of a progenitor of a Milky Way sized galaxy selected from the sample of MW progenitors from the Illustris TNG50 simulation (Pillepich et al., 2018; Nelson et al., 2019a) with an early forming disk resembling the chemo-kinematic structure of the MW stellar disk (Semenov et al., 2024a). The galaxy was re-simulated using the ART code at higher resolution and including a detailed modeling of multiphase ISM, radiative and stellar feedback, subgrid turbulence, and a model of variable star formation efficiency (Kravtsov, 1999; Rudd et al., 2008; Gnedin & Kravtsov, 2011; Semenov et al., 2016, 2018). The model details are described in Semenov et al. (2024a), and we refer to this simulation as S24 ART. No IGM gas metallicity floor was assumed in this model.
Semenov et al. (2024b) present a detailed analysis of the early evolution of the MW progenitor in this simulation and show that at $z\gtrsim 6-7$ the progenitor has a very bursty star formation and drives significant outflows. In this work, we study the metallicity distribution of the IGM gas accretable by UFD progenitors within the zoom-in region of this simulation at redshifts $5\leq z\leq 10$.
@@ 2.2.3 MEGATRON simulations
At the current time, only the snapshots at $z\gtrsim 8.5$ are available and are used in our analyses. As noted above, for the FIRE-2 and S24 ART simulations, due to truncation of star formation due to reionization, the IGM metallicity distribution at $z\gtrsim 8$ has the greatest impact on the metallicity distribution of stars in model UFD galaxies.
@@ 3.1 Enrichment of the IGM gas and its impact on the stellar metallicity of UFDs
We first consider the metallicity of the IGM that would be accretable by UFDs in a MW-like environment at redshifts $105<z<$ 10 where they are expected to be star-forming. The goal is to gauge whether the enrichment of accretable IGM gas that is predicted by hydrodynamical simulations reproduces the mean stellar metallicity and the metallicity scatter of observed UFDs around the Milky Way.
To this end, we model the evolution of UFDs using the GRUMPY fiducial model and Caterpillar MAHs, but assuming models with a probability distribution function (pdf) of the accreted IGM metallicity, $Z_{\rm IGM}$, instead of the constant $Z_{\rm IGM}$ value of the fiducial model. We adopt the pdfs for IGM metallicity based on the gas that is accretable by UFD progenitors in several cosmological hydrodynamic simulations of the MW progenitor galaxies at $z\gtrsim 5$.
@@ 3.1.1 Defining the gas accretable by UFDs
To determine the metallicity of gas accretable by UFDs across redshifts, we first need to identify gas that UFD progenitors can potentially accrete. To this end, we use Caterpillar evolutionary tracks of model UFDs, defined as galaxies with $M_{V}>-7.7$ at $z=0$, following the UFD limit adopted by Simon (2019). We also limit analysis to the dwarfs that have apparent $V$-band magnitudes of $m_{V}<17.5$ if observed from the center of their parent halo at $z=0$, which approximately corresponds to the apparent magnitudes of UFDs identified in SDSS, DES, and PanSTARRS surveys (see Drlica-Wagner et al., 2020). This limits the model galaxies to the luminosities of observed galaxies sufficiently bright for identification and metallicity measurements. We also restrict galaxies to a distance of $r\leq 100\ \rm kpc$ from the center of the host halo at $z=0$. This was done originally to alleviate concerns that the analyzed Lagrangian regions in some cosmological simulations are small. However, we found that the size of the analyzed region has little effect on the distributions of UFD progenitor galaxy properties at the redshifts we sample. We use the tracks of the model galaxy sample defined in this way to identify locations of their progenitors at $105<z<$ 10 relative to the main progenitor of the host MW-sized galaxy and their virial masses.
@@
We use distances to the host progenitor, $M_{200}$, and $T_{200}$ values sampled from the distributions shown in Figure 1 and gas properties in galaxy formation simulations to identify gas accretable by UFD progenitors. Specifically, we consider a gas parcel to be accretable by a UFD if 1) its density is below $100\ \overline{\rho}_{\rm b}(z)$, thereby excluding gas within bound structures, 2) its temperature is lower than $T_{200}$, and 3) it is located within $R_{200}$ of the mock UFD progenitor. We then calculate the mass-averaged metallicity distribution of the accretable gas by repeating the process for tens of thousands of mock halo locations and properties at each redshift.
@@ 3.1.2 Metallicity distribution of the IGM accretable by UFDs
Figure 2: Cumulative distributions of the metallicity of gas available for accretion by UFD progenitor halos in the environment around simulated MW analogs at several redshifts. Average stellar metallicities in the GRUMPY model are taken to be defined by the mass of metals and total mass in stars relative to solar. Note that the axis limits vary from panel to panel, however in all model families, only a small fraction of enriched gas is available at the highest and most impactful redshifts. Left: All 8 FIRE-2 m12 simulations, where model-to-model variation in gas availability is sufficiently low that the sample data from all models are combined to create this distribution. These models impose a metallicity floor of $\mathrm{[Fe/H]}\simeq-3.9$. Center: The S24 ART model. Unlike FIRE-2 models, no metallicity floor imposed in the model. The lower [Fe/H] limit of this figure is selected for convenience of visualization to show the distribution in a region below $\mathrm{[Fe/H]}=-4$, however the majority of available gas resides at significantly lower metallicities. Proportionally more gas than for other MW-like models is enriched above $\mathrm{[Fe/H]}=-2$ as this model evolves, however at $z=10$ it is the least enriched. Right: Several MEGATRON models with different feedback or star formation prescriptions are shown with different line styles. Distributions for $z=10$ and $z=8.5$ are shown with the same color scheme as FIRE-2 legend. Lower redshift data is not yet available at the time of writing. Unlike other models shown, MEGATRON models include explicit Pop III stellar prescriptions in addition to no imposed metallicity floor. In general, increasing feedback strength or changing the IMF leads to more available metals in gas near the host galaxy progenitor compared to the fiducial (Efficient SF) model. Similar to the center panel, the majority of available gas resides at significantly lower metallicity than shown.
@@
The middle panel of Figure 2 shows the metallicity distribution in the S24 ART model of an early forming MW analog. The plot shows that in the S24 simulation 98% of the gas available to UFDs at $z>7$ has $\rm[Fe/H]<-4$. At $z\approx 5-7$, however, the amount of enriched gas increases rapidly with $\approx 10\%$ of the accretable gas at $\rm[Fe/H]>-4$ at $z=5$. In addition, the range of metallicity of accretable gas in this simulation is wider than in the FIRE-2 simulations and there is a more prominent tail of gas extending to metallicities $\rm[Fe/H]>-2$ than in other simulations.
@@
The right panel of Figure 2 shows cumulative distributions of $Z_{\rm IGM}$ at $z\geq 8.5$ in several MEGATRON re-simulations of a progenitor of MW-sized galaxy with varying feedback and star formation prescriptions, including explicit modeling of enrichment due to Pop III stars that are not included in the FIRE-2 and S24 simulations. These simulations also do not assume any minimum value of the gas metallicity.
@@
Of the MEGATRON simulations analyzed, the gas accretable by UFDs has the lowest metallicities in the efficient SF simulations. The gas has the broadest range of metallicities in the bursty SF model, and therefore the greatest proportional availability of higher metallicity gas, reflecting the stronger outflows in this simulation. The width of the metallicity distribution in the variable IMF model is intermediate between the efficient SF and bursty SF simulations. The type II supernova energy in this simulation is the same as in the efficient SF run, but a metallicity and density-dependent IMF that becomes top-heavy in some region, and inclusion of energy injection from hypernovae, lead to increased enrichment. The differences between these models is minor, however, and in our analyses below we consider only the bursty SF model.
@@
All of the simulations examined here have strong feedback-driven outflows and/or an early MW progenitor formation time, as explicitly chosen in MEGATRON and S24 simulations. In addition to strong feedback ejecting more metals from the host, early forming galaxies have had more time to evolve and enrich their environments by a given redshift. Therefore the level of gas enrichment in these simulations may be considered a conservative upper limit on what can be expected for the Milky Way environments at $z>5$. Nevertheless, as we show below, this level of enrichment has a negligible effect on the metallicity distribution of stars in UFDs in the GRUMPY model.

### v2

@@ 1 Introduction
Currently, there is no established physical interpretation of the metallicities of UFD galaxies and their scatter in terms of the physical processes shaping the evolution of these galaxies. In this study, we attempt to construct such an interpretation using a combination of the IGM metallicity distributions derived from several state-of-the-art cosmological simulations of galaxy formation and a semi-analytical model of galaxy formation based on a large number of mass assembly histories of UFD progenitors residing in the Milky Way-sized halos at $z=0$. Specifically, we explore a range of models with different assumptions in modeling the accreted IGM metallicity and different outflow mass loading factors. We compare the luminosity–metallicity relation of the model galaxies and their stellar metallicity distribution functions (MDFs) to existing measurements for observed UFDs with the goal of gauging the relative importance of external IGM enrichment and outflows in setting UFD metallicities.
@@ 2.1 Model of UFD evolution
Our model for the evolution of ultra-faint galaxies is based on the halo evolution tracks extracted from the Caterpillar [^1] https://www.caterpillarproject.org suite of cosmological zoom-in $N$-body simulations of Milky Way-mass ($0.5\lesssim M/10^{12}\,M_{\odot}\lesssim 2$) halos (Griffen et al., 2016). Specifically, we use the $32$ LX14 Caterpillar simulations with the dark matter particle mass of $3\times 10^{4}M_{\odot}$. This resolution has been shown to be sufficient to reliably model subhalos down to peak masses (largest mass during halo evolution) of $\gtrsim 10^{7}\,M_{\odot}$ (Pham et al., 2023). The objects we model in the current study form in halos of $\geq 10^{8}\,M_{\odot}$, well above this limit. Combining these suites provides a population of several thousand halos in MW-like environments, which we use in our study.
@@
[…]
@@
In cosmological simulations such as FIRE-1 (Muratov et al., 2015), FIRE-2 (Pandya et al., 2021), EAGLE (Mitchell et al., 2020), and Illustris TNG50 (Nelson et al., 2019b), a clear trend of increasing average $\eta$ with decreasing stellar mass is present. However, these analyses do not probe galaxies with stellar mass as low as UFDs $(M_{\star}\lesssim 10^{5.5}M_{\odot})$, so the extent to which this relation can be extrapolated to galaxies of stellar mass $M_{\star}<10^{6}\,M_{\odot}$ is unclear. Furthermore, the mass loading factor of galaxies at high redshifts where UFDs are expected to form most of their stars is also largely unknown.
@@
The fiducial values of Manwadkar & Kravtsov (2022) adopted in this work are $\eta_{\rm norm}=1.8$, $\eta_{p}=-0.45$, $\eta_{c}=4.5$, $M_{\star,10}\equiv M_{\star}/10^{10}M_{\odot}$, and $\eta_{\rm max}=2000$. These values were chosen to be close to those in the FIRE-1 and FIRE-2 simulations for galaxies of $10^{6}\lesssim M_{\star}/M_{\odot}\lesssim 10^{9}$ (Muratov et al., 2015; Pandya et al., 2021), while at smaller and larger masses the power law relation $\eta\propto M_{\star,10}^{\eta_{p}}$ is modified to reproduce the observed stellar mass to halo mass relation (SHMR) and stellar mass–metallicity relations, as detailed in Manwadkar & Kravtsov (2022). The subtraction of $\eta_{\rm c}$ results in the rapid decrease of $\eta$ for $M_{\star}\gtrsim 10^{10}\,M_{\odot}$, while the maximum value of $\eta_{\rm max}=2000$ was chosen to account for the expectation in the UFD regime the decreasing stellar mass-to-halo mass ratio results in less efficient driving in outflows and to prevent unlimited increase of $\eta$ at the very low stellar masses.
The values of other parameters are set to the fiducial values adopted in Manwadkar & Kravtsov (2022): $y_{Z}=0.06$ for the Chabrier (2003) IMF (e.g., Vincenzo et al., 2016), $\mathcal{R}=0.34$, $\zeta_{\rm w}=1$. The expected value and scaling of $\zeta_{\rm w}$ with system mass are currently not known. Manwadkar & Kravtsov (2022) showed that a good match to the observed luminosity–metallicity relation of bright dwarf galaxies can be obtained with $\zeta_{\rm w}=1$, as long as a mass-dependent scaling of $\eta$ is adopted. Results of the FIRE-1 and FIRE-2 galaxy formation simulations are also consistent with $\zeta_{\rm w}\approx 1$ independent of galaxy mass (Muratov et al., 2017; Pandya et al., 2021), albeit with a sizable scatter.
@@ 2.2 Galaxy formation simulations
We consider the properties of these simulated MW-like environments in the redshift range $5<z<10$, corresponding to the redshift range of the publicly available FIRE-2 high redshift suite. We found that the metallicity of the IGM at $z\approx 8-10$ has the most effect on the stellar metallicities of model UFDs within this range (see Section 3.1). We use the $z=10$ metallicity distribution to model IGM metallicity in the GRUMPY model at $z>10$ – an overestimate of the metallicity of accreted gas, but a conservative choice that strengthens our conclusions below. Similarly, we extend the $z=5$ metallicity distribution to $z=0$ in the GRUMPY model, where the choice of $Z_{\rm IGM}$ has little impact on model UFD metallicities because UFDs are no longer star forming.
@@ 2.2.1 FIRE-2 simulations
We consider simulation outputs in the redshift range $5<z<10$, the range publicly available in both the core m12 models (Wetzel et al., 2016; Garrison-Kimmel et al., 2017, 2019; Hopkins et al., 2018; Samuel et al., 2020) and the high redshift simulations of smaller-mass halos (e.g., Ma et al., 2018, 2019, 2020), suitable to probe the environments of UFDs prior to cosmic reionization when most of their stars are expected to form. All FIRE-2 simulations analyzed here assume an IGM gas metallicity floor of $\mathrm{[Fe/H]}\simeq-3.9$.
2.2.2 Semenov et al. 2025 simulation of a Milky Way progenitor
We also study a simulation of a progenitor of a Milky Way sized galaxy selected from the sample of MW progenitors from the Illustris TNG50 simulation (Pillepich et al., 2018; Nelson et al., 2019a) with an early forming disk resembling the chemo-kinematic structure of the MW stellar disk (Semenov et al., 2025b). The galaxy was re-simulated using the ART code at higher resolution and including a detailed modeling of multiphase ISM, radiative and stellar feedback, subgrid turbulence, and a model of variable star formation efficiency (Kravtsov, 1999; Rudd et al., 2008; Gnedin & Kravtsov, 2011; Semenov et al., 2016, 2018). The model details are described in Semenov et al. (2025b), and we refer to this simulation as S25 ART. No IGM gas metallicity floor was assumed in this model.
Semenov et al. (2025a) present a detailed analysis of the early evolution of the MW progenitor in this simulation and show that at $z\gtrsim 6-7$ the progenitor has a very bursty star formation and drives significant outflows. In this work, we study the metallicity distribution of the IGM gas accretable by UFD progenitors within the zoom-in region of this simulation at redshifts $5\leq z\leq 10$.
@@ 2.2.3 MEGATRON simulations
At the current time, only the snapshots at $z\gtrsim 8.5$ are available and are used in our analyses. As noted above, for the FIRE-2 and S25 ART simulations, due to truncation of star formation due to reionization, the IGM metallicity distribution at $z\gtrsim 8$ has the greatest impact on the metallicity distribution of stars in model UFD galaxies.
@@ 3.1 Enrichment of the IGM gas and its impact on the stellar metallicity of UFDs
We first consider the metallicity of the IGM that would be accretable by UFDs in a MW-like environment at redshifts $5<z<10$ where they are expected to be star-forming. The goal is to gauge whether the enrichment of accretable IGM gas that is predicted by hydrodynamical simulations reproduces the mean stellar metallicity and the metallicity scatter of observed UFDs around the Milky Way.
To this end, we model the evolution of UFDs using the GRUMPY fiducial model and mass assembly histories of the Caterpillar halos, but assuming models with a probability distribution function (pdf) of the accreted IGM metallicity, $Z_{\rm IGM}$, instead of the constant $Z_{\rm IGM}$ value of the fiducial model. We adopt the pdfs for IGM metallicity based on the gas that is accretable by UFD progenitors in several cosmological hydrodynamic simulations of the MW progenitor galaxies at $z\gtrsim 5$.
@@ 3.1.1 Defining the gas accretable by UFDs
To determine the metallicity of gas accretable by UFDs across redshifts, we first need to identify gas that UFD progenitors can potentially accrete. To this end, we use Caterpillar evolutionary tracks of model UFDs, defined as galaxies with $M_{V}>-7.7$ at $z=0$, following the UFD limit adopted by Simon (2019). We also limit analysis to the dwarfs that have apparent $V$-band magnitudes of $m_{V}<17.5$ if observed from the center of their parent halo at $z=0$, which approximately corresponds to the apparent magnitudes of UFDs identified in SDSS, DES, and PanSTARRS surveys (see Drlica-Wagner et al., 2020). This limits the model galaxies to the luminosities of observed galaxies sufficiently bright for identification and metallicity measurements. We also restrict galaxies to a distance of $r\leq 100\ \rm kpc$ from the center of the host halo at $z=0$. This was done originally to alleviate concerns that the analyzed Lagrangian regions in some cosmological simulations are small and also because most observed UFDs are located within 100 kpc of Milky Way. However, we found that the size of the analyzed region has little effect on the distributions of UFD progenitor galaxy properties at the redshifts we sample. Furthermore, the closest surviving objects tend to be at smaller distances to the progenitor of the host galaxy at high $z$, which are generally more enriched. This makes the main conclusion of our study conservative.
We use the tracks of the model galaxy sample defined in this way to identify locations of their progenitors at $5<z<10$ relative to the main progenitor of the host MW-sized galaxy and their virial masses.
@@
[…]
As we noted in Section 2.1, Pham et al. (2023) showed that the LX14 Caterpillar simulations reliably model subhalos down to peak masses of $\gtrsim 10^{7}\,M_{\odot}$. Although some of the objects in Figure 1 have smaller masses at high $z$, their peak masses are $10^{8}\,M_{\odot}$ and are well inside the regime reliably resolved in the simulations.
@@ 3.1.2 Metallicity distribution of the IGM accretable by UFDs
Figure 2: Cumulative distributions of the metallicity of gas available for accretion by UFD progenitor halos in the environment around simulated MW analogs at several redshifts. Average stellar metallicities in the GRUMPY model are taken to be defined by the mass of metals and total mass in stars relative to solar. Note that the axis limits vary from panel to panel, however in all model families, only a small fraction of enriched gas is available at the highest and most impactful redshifts. Left: All 8 FIRE-2 m12 simulations, where model-to-model variation in gas availability is sufficiently low that the sample data from all models are combined to create this distribution. These models impose a metallicity floor of $\mathrm{[Fe/H]}\simeq-3.9$. Center: The S25 ART model. Unlike FIRE-2 models, no metallicity floor imposed in the model. The lower [Fe/H] limit of this figure is selected for convenience of visualization to show the distribution in a region below $\mathrm{[Fe/H]}=-4$, however the majority of available gas resides at significantly lower metallicities. Proportionally more gas than for other MW-like models is enriched above $\mathrm{[Fe/H]}=-2$ as this model evolves, however at $z=10$ it is the least enriched. Right: Several MEGATRON models with different feedback or star formation prescriptions are shown with different line styles. Distributions for $z=10$ and $z=8.5$ are shown with the same color scheme as FIRE-2 legend. Lower redshift data is not yet available at the time of writing. Unlike other models shown, MEGATRON models include explicit Pop III stellar prescriptions in addition to no imposed metallicity floor. In general, increasing feedback strength or changing the IMF leads to more available metals in gas near the host galaxy progenitor compared to the fiducial (Efficient SF) model. Similar to the center panel, the majority of available gas resides at significantly lower metallicity than shown.
@@
The middle panel of Figure 2 shows the metallicity distribution in the S25 ART model of an early forming MW analog. The plot shows that in the S25 simulation 98% of the gas available to UFDs at $z>7$ has $\rm[Fe/H]<-4$. At $z\approx 5-7$, however, the amount of enriched gas increases rapidly with $\approx 10\%$ of the accretable gas at $\rm[Fe/H]>-4$ at $z=5$. In addition, the range of metallicity of accretable gas in this simulation is wider than in the FIRE-2 simulations and there is a more prominent tail of gas extending to metallicities $\rm[Fe/H]>-2$ than in other simulations.
@@
The right panel of Figure 2 shows cumulative distributions of $Z_{\rm IGM}$ at $z\geq 8.5$ in several MEGATRON re-simulations of a progenitor of MW-sized galaxy with varying feedback and star formation prescriptions, including explicit modeling of enrichment due to Pop III stars that are not included in the FIRE-2 and S25 simulations. These simulations also do not assume any minimum value of the gas metallicity.
@@
Of the MEGATRON simulations analyzed, the gas accretable by UFDs has the lowest metallicities in the efficient SF simulations. The gas has the broadest range of metallicities in the bursty SF model, and therefore the greatest proportional availability of higher metallicity gas, reflecting the stronger outflows in this simulation. The width of the metallicity distribution in the variable IMF model is intermediate between the efficient SF and bursty SF simulations. The type II supernova energy in this simulation is the same as in the efficient SF run, but a metallicity and density-dependent IMF that becomes top-heavy in some regime, and inclusion of energy injection from hypernovae, lead to increased enrichment. The differences between these models is minor, however, and in our analyses below we consider only the bursty SF model.
@@
All of the simulations examined here have strong feedback-driven outflows and/or an early MW progenitor formation time, as explicitly chosen in MEGATRON and S25 simulations. In addition to strong feedback ejecting more metals from the host, early forming galaxies have had more time to evolve and enrich their environments by a given redshift. Therefore the level of gas enrichment in these simulations may be considered a conservative upper limit on what can be expected for the Milky Way environments at $z>5$. Nevertheless, as we show below, this level of enrichment has a negligible effect on the metallicity distribution of stars in UFDs in the GRUMPY model.